# Universal quantum control over bosonic network

### v1

@@ V Chiral state transfer in network
In addition, the preceding passage along $\mu_{4}(t)$ can be straightforwardly extended to run a four-mode chiral transfer by a four-stage passage, i.e., in Stage (i) for $t\in[0,\tau]$, $a_{1}\rightarrow a_{2}$; in Stage (ii) for $t\in[\tau,2\tau]$, $a_{2}\rightarrow a_{3}$; in Stage (iii) for $t\in[2\tau,3\tau]$, $a_{3}\rightarrow a_{4}$; and in Stage (iv) for $t\in[3\tau,4\tau]$, $a_{4}\rightarrow a_{1}$. The boundary conditions for $\theta_{k}(t)$ with $1\leq k\leq 3$ of Stages (i) and (ii) remain invariant as those for the three-mode transfer protocol. And in Stages (iii) and (iv), they are set as $\theta_{3}(3\tau)=\pi$ and $\theta_{1}(4\tau)=\theta_{2}(4\tau)=\theta_{3}(4\tau)=\pi/2$, respectively. In practical, $\theta_{1}(t)$ and $\theta_{2}(t)$ in Stages (i-ii) and $\theta_{3}(t)$ in Stages (i-ii) and (iv) can chosen the same as Eqs. (54) and (55), respectively. And one can choose $\theta_{1}(t)=\theta_{2}(t)=\Phi(t)+\pi/2$ for Stages (iii-iv) and $\theta_{3}(t)=(\pi/2)[1-\sin(\pi t/(2\tau))]$ for Stage (iii).
VI Conclusion
@@
In summary, we propose a theoretical framework to construct nonadiabatic passages for the general $N$-mode bosonic network that are governed by the time-dependent Hamiltonian. With a completed set of time-independent ancillary modes, we find a necessary and sufficient condition to exactly solve the time-dependent Schrödinger equation or equivalently to determine the dynamics of ancillary operators in the Heisenberg picture. In particular, the diagonalization of the coefficient matrix for Hamiltonian and gauge potential in the representation of the time-independent ancillary modes can be implemented by its commutation with the projection operators. The commutation condition yields the parametric constraints for the network Hamiltonian, which activate the ancillary modes as versatile nonadiabatic passages. Along the activated passages, arbitrary states can be exchanged between any pair of modes in the bosonic network. As illustrative examples, the feasibility of our theory is confirmed by perfect state-exchange in a two-mode network, the chiral NOON-state transfer in a three-mode system, and the chiral Fock-state transfer among three of four bosonic modes. Our work thus provides a universal approach for controlling quantum bosonic networks with arbitrary full connection and size.
@@ Acknowledgments
We acknowledge grant support from the Science and Technology Program of Zhejiang Province (No. 2025C01028).
@@
$-\sin\theta_{k+1}(0)e^{-i\frac{\alpha_{k+1}(0)}{2}}a_{k+2},$

### v2

@@ V Chiral state transfer in network
Moreover, the ideal constraint equations (53) and (LABEL:CondDetu) can be slightly modified to run a four-mode chiral transfer by a four-stage passage along the ancillary mode $\mu_{4}(t)$. In particular, the control target becomes $a_{1}\rightarrow a_{2}$ in Stage (i) for $t\in[0,\tau]$, $a_{2}\rightarrow a_{3}$ in Stage (ii) for $t\in[\tau,2\tau]$, $a_{3}\rightarrow a_{4}$ in Stage (iii) for $t\in[2\tau,3\tau]$, and $a_{4}\rightarrow a_{1}$ in Stage (iv) for $t\in[3\tau,4\tau]$. For this target, the boundary conditions for $\theta_{k}(t)$ with $1\leq k\leq 3$ of Stages (i) and (ii) remain invariant as those for the three-mode transfer protocol. And in Stages (iii) and (iv), they are reset as $\theta_{3}(3\tau)=\pi$ and $\theta_{1}(4\tau)=\theta_{2}(4\tau)=\theta_{3}(4\tau)=\pi/2$, respectively. In practice, $\theta_{1}(t)$ and $\theta_{2}(t)$ in Stages (i-ii) and $\theta_{3}(t)$ in Stages (i-ii) and (iv) can be the same as in Eqs. (55) and (56), respectively. And one can choose $\theta_{1}(t)=\theta_{2}(t)=\Phi(t)+\pi/2$ for Stages (iii-iv) and $\theta_{3}(t)=(\pi/2)[1-\sin(\pi t/(2\tau))]$ for Stage (iii).
VI Discussion and Conclusion
@@
In summary, we propose a theoretical framework to construct nonadiabatic passages for the general $N$-mode bosonic network that is governed by the time-dependent Hamiltonian. With a completed set of time-independent ancillary modes, we find a necessary and sufficient condition to exactly solve the time-dependent Schrödinger equation or equivalently to determine the dynamics of ancillary operators in the Heisenberg picture. In particular, the diagonalization of the coefficient matrix for Hamiltonian and gauge potential in the representation of the time-independent ancillary modes can be implemented by its commutation with the projection operators of the matrix bases. The commutation condition yields the parametric constraints for the network Hamiltonian, which activate the ancillary modes as versatile nonadiabatic passages. Along the activated passages, arbitrary states can be exchanged between any pair of modes in the bosonic network, which is robust against the unknown crosstalk among relevant nodes, the environmental decoherence, and the imperfect parametric setting. As illustrative examples, the feasibility of our theory is confirmed by the perfect state exchange in a two-mode network, the chiral NOON-state transfer in a three-mode system, and the chiral Fock-state transfer among three of four bosonic modes. Our work thus provides a universal approach for controlling quantum bosonic networks with arbitrary connection and size.
Our protocol for the bosonic network connected with beam-splitter coupling can be realized in a variety of physical platforms. They include and are not limited to the circuit quantum electrodynamics system Vermersch et al. (2017); Blais et al. (2021), which is about two cavity modes; the circuit quantum acoustodynamics system von Lüpke et al. (2024), in which a mechanical mode is simultaneously coupled to two other mechanical modes; the cavity magnomechanical systems Zhang et al. (2016a); Li et al. (2021); Shen et al. (2022, 2025), in which a cavity-magnon polariton is coupled to a phonon mode; and the cavity magnonic systems Zhang et al. (2014, 2016b); Lachance-Quirion et al. (2019); Xu et al. (2021); Zare Rameshti et al. (2022); Xu et al. (2023); Zheng et al. (2023), where a cavity mode is coupled to a magnon mode. In addition, the cavity magnonic systems can be directly extended to the multi-mode case Zhang et al. (2015b), where multiple magnon modes are coupled to a common cavity mode. In comparison to the existing methods, our protocol is free of at least three typical limitations: (1) the running time of protocol, (2) the capability to transfer arbitrary states, and (3) the scalability to a large size.
@@ Acknowledgments
We acknowledge grant support from the National Natural Science Foundation of China (Grant No. U25A20199) and the “Pioneer” and “Leading Goose” R&D of Zhejiang Province (Grant No. 2025C01028).
@@
$-\sin\theta_{k+1}(0)e^{-i\frac{\alpha_{k+1}(0)}{2}}a_{k+2}=\mu_{k+1}(0),$